OpenStudy (anonymous):

solve the system of elmination method 5x+3y=-7 7x-2y=11

OpenStudy (anonymous):

multiply the first equation by 2, the second by 3, then add

OpenStudy (anonymous):

\[10x+6y=-14\] \[21x-6y=33\] \[31x=19\] \[x=\frac{19}{31}\]

OpenStudy (anonymous):

solve for x or y in either equation. let's use the first. 3y=-7 - 5x; thus, y = -7/3 -5/3x. plug this into the second equation to get: 7x-2(-7/3-5/3x)=11, expand to get 7x+14/3+10/3x=11, simplifly and reduce to get x= 19/31, then just plug this back into the first equation to find the value of y.

OpenStudy (anonymous):

replace x by \[\frac{19}{31}\] in either equation to find y

OpenStudy (anonymous):

@tsizzle that method will of course work, but question said use "elimination"

OpenStudy (anonymous):

you will get \[y=-\frac{104}{31}\]

OpenStudy (anonymous):

yeah, my bad, wasn't really paying that much attention :/ haha, I'm watching tv at the same time and just solving pellet